OpenStudy (anonymous):

Let A, B, C denote three rectangles, all with perimeter equal to 240, whose heights are h_A = 20, h_B = 60, h_C = 110, respectively. Which of the rectangles A, B, or C has the smallest area?

OpenStudy (anonymous):

for A length =240/2 -20 =100 Area =100*20 =2000 for B length =240/2 - 60 =60 Area =60*60 =3600 for C length =240/2 -110=10 Area =10*110 =1100 C has the smallest area